# Exploring Young Stellar Variability with Gaia DR3 light curves

### v1

@@ 6.1 Incidence of morphological classes across surveys
We also point to survey-to-survey differences in the time span of observations. High-cadence surveys cover between 40 and 80 days of observations, while ZTF surveys cover 800-1 500 days, and GDR3 750-1 000 days. These differences imply that quasiperiodic phenomena at intermediate timescales may be perceived as aperiodic by studies with reduced time span. Moreover, L-variables are defined as sources with timescales larger than 20-40 d in high-cadence studies, 100-250 d in ZTF, we use 350-500 d in the present study. L-variable fractions are thus fully survey-dependent. Even if removed from our class incidence comparison, their impact on the definition of other morphological classes is still worth mentioning. First, L-variables could be operationally defined as aperiodic, since their time scale was not resolvable over the duration of a survey. Although their lack of reliable timescale implies that their Q indices cannot be estimated, their M indices reveal that they would increase class frequencies by 3.8% for B, 8.6% for S, and 5.3% for APDs - thus bringing our class incidence for S and APD variables to levels similar to those of ZTF studies. Second, the timescale limit for the L-variable definition indirectly determines how many folds are used in the waveform estimation required to derive the Q index. With our threshold of half of the total duration of observations, only one fold is included for sources with the longest timescales, which may lead to over-fitting of estimated waveforms, resulting in deflated Q-indices, and causing aperiodic phenomena to be miss-detected as quasiperiodic.
@@
Some previous studies attributed multiple morphological classes to multi-mode variables (e.g. Cody & Hillenbrand, 2018; Cody et al., 2022; Hillenbrand et al., 2022) also including an additional class, multiperiodic, to account for variables with multiple periods. To improve the comparison of class frequencies with previous studies, in Table 4 we removed sources classified as multiperiodic and adopted the primary morphological class when multiple are reported. Nevertheless, with our focus on the dominant mode of variability, we acknowledge that undetected multi-mode variables in our sample may display inflated Q-indices. Beyond multiple variability mechanisms that occur simultaneously, the longer timespan of surveys like Gaia and ZTF may also facilitate the observation of sources with alternating dominant variability phenomena. For example, this can result in symmetric M-indices derived from light curves that shift from bursting to dipping behaviour (an example is shown in Fig. 20).
@@
Our combined fraction of dipping variables (QPD+APD) is equivalent to the one found in ZTF studies, but smaller than the one found in high-cadence studies. Nevertheless, as already noted, our APD fractions are underestimated due to our definition of L-variables. Furthermore, in Sects. 6.2.3 and 6.2.4 we discuss evidence that our sample of dippers includes a different underlying population of YSOs in comparison to the high-cadence surveys. Similarly, B-variables also include different variability mechanisms, and in Sect. 6.2.5 we discuss evidence that Gaia’s sampling yields the detection of a distinct population of bursting variables in comparison with other surveys.
6.2 General properties of Morphological Classes
@@ 6.2.1 Strictly Periodic
Strictly periodic variability, P (see Fig. 15), is typically associated with rotational modulation by magnetically induced cool spots at the stellar surface in sources viewed close to edge-on (; Cody et al., 2022). From the top-left panel of Fig. 14, the timescale distribution for P sources is consistent across studies and dominated by rotational timescales. The variability amplitude in our primary sample (blue) is higher than that reported in high-cadence surveys (red) but slightly lower than in ZTF studies (orange); the secondary sample (grey) exhibits amplitude levels comparable to those of high-cadence studies. Differences in amplitude distributions can be attributed to selection effects in different surveys. Typically, cool spots have lifetimes ranging from 1 to 50 rotation cycles (Basri et al., 2022). Higher-cadence surveys, which usually span 40 to 80 d, provide excellent phase coverage before significant changes in spot coverage occur. In Gaia, sparse sampling and long-term coverage implies that hundreds of rotation cycles take place before achieving sufficient phase coverage, thus biasing the sample of P sources towards sources with larger and more stable spot coverage (and large amplitudes). In contrast, the higher noise level of ZTF’s ground-based observations allows only the higher-amplitude variables to be detected. As a result, Fig. 14 shows that the ZTF distributions of all classes are shifted toward higher amplitudes compared to other surveys.
The properties of our sample of P-variables support the hypothesis that their variability is primarily driven by rotational modulation linked to relatively stable cool spots. Their average amplitude is among the lowest of all morphological classes, and the ratio between rotation and intermediate timescales (Table 3) shows that this sample is dominated by rotational modulation. P-variables are 82.95% composed of YSOs without evidence of a thick disk or envelope. Furthermore, 86.27% of P-sources with H information are WTTS.
@@ 6.2.2 Quasi-periodic Symmetric
As discussed in Sec. 6.1, QPS (see Fig. 16) is the most common morphological class among our entire sample. Multiple physical origins have been proposed for QPS. Although generally linked to rotational modulation of stellar spots, high-amplitude QPS have been tentatively associated with variability caused by large hot spots induced by accretion, in combination with underlying lower-amplitude accretion variation (Venuti et al., 2015; Cody & Hillenbrand, 2018), while low-amplitude QPS are thought to originate from changes of cool spot coverage due to magnetic activity (Cody et al., 2022). Considering that cool spots are the only variability driver expected in diskless sources and that disk-bearing and actively accreting sources are likely to exhibit accretion-induced hot spots, this view is generally supported by our sample. We find that the average distribution of amplitudes is lower for diskless and thin-disk QPS (0.11 mag) compared to sources with envelope or thick disk (0.30). Similar differences are found when comparing the amplitudes of the WTTS QPS (0.11) with the CTTS QPS (0.36). The same trend is verified by comparing our primary and secondary samples, where the CTTS frequency varies from 36.6% to 6.1%.
@@
Our QPS-variables include a population of sources with intermediate timescales, which has not been found in previous studies. Fig. 14 shows that for the primary sample, the timescale distribution is bimodal, with 1.3 times as many rotation timescales as intermediate ones. A comparison of their Q-index distributions reveals significant differences (with KS test support), where the rotational-timescale QPS are shifted toward higher periodicity with an average Q of 0.34 compared to 0.44 for intermediate-timescale QPS. Visual inspection of intermediate-timescale QPS light curves revealed that many of these sources have limited phase coverage when folded to their characteristic timescale. This limitation is inherent in surveys with sparse sampling as Gaia, where some phase-folded light curves with poor phase coverage include large gaps with no data. These gaps, by turn, bring us to the limits of the methodology for waveform estimation, causing waveforms to be over-fit to only a fraction of the phase, artificially shifting Q-indices towards smaller values and causing miss-detection of aperiodic sources as quasi-periodic. With Gaia’s scanning law, the timescales that yield gaps in phase coverage vary with sky position, making modelling and masking these timescales from our analysis unpractical. To mitigate this issue, further studies should consider the complementation of Q&M-analysis with an additional variability metrics sensitive to the degree of phase coverage as part of the procedure for the derivation of the Q index (e.g. Saunders et al., 2006). We also considered adjusting the quasi-periodic/aperiodic Q-index threshold to mitigate this effect, but verified that a lower Q-index limit would result in reliable quasi-periodic light-curves with rotation timescales being misclassified as aperiodic.
@@ 6.2.3 Quasi-periodic Dippers
QPD light curves (see Fig. 17) are often associated with the phenomenology of AA Tauri variables (Bouvier et al., 1999). In such cases, a misalignment between the rotation and magnetic field axes yields the formation of wraps, with disk material lifted above the disk’s mid-plane and close to the co-rotation radius, introducing quasi-periodic occultations at rotational timescales in systems observed close to edge-on (Romanova et al., 2013; McGinnis et al., 2015). Among the primary sample, 62.09% of QPD have an inner disk, and 40.60% have evidence of ongoing accretion (CTTS classification), supporting the association of this class with physical phenomenology related to the presence of a disk and ongoing accretion. However, a direct association of these with the AA Tau prototype is hampered by two factors.
@@
Firstly, we noted that a fraction of our low-amplitude QPDs may be misclassified QPS light-curves. About half of our QPDs are in the secondary sample, which shows a much lower disk frequency (13.71%), and even lower CTTS frequency (4.47%), with both Q and M distributions shifted in relation to the primary sample towards more symmetric M-indices and more periodic Q-indices. Fig. 14 shows that these have systematically lower amplitudes than QPD from high-cadence surveys. The M-index distribution for the secondary QPD sample peaks close to the threshold for dipper selection with 90% of the M-indices for this sample below the median value for the primary QPD sample. Additionally, although statistically different in shape, the distributions of amplitudes and Q-index for the secondary QPD and secondary QPS samples have similar averages and standard deviations, as well as similar disk and CTTS frequency. Finally, visual examination of the light curves also suggested that a higher M-index threshold might be necessary to enhance the purity of the dipper variable samples, where a reviewed threshold would identify a large fraction of secondary QPD as QPS.
@@
Secondly, because of Gaia’s much longer observation time span, QPDs may be sensitive to other dimming variables beyond AA Tau. Fig. 14 reveals a bimodal distribution for QPDs, with of QPDs showing intermediate timescales. This contradicts the predictions of the AA Tau phenomenology, where the occulting material is anticipated to be near the co-rotation radius. The primary sample of QPDs with intermediate timescales shows high disk and CTTS frequencies (65.09, and 47.35%), suggesting a physical origin also associated with the presence of a disk and ongoing accretion. Statistical tests for distribution shape, mean and variance differences for the distributions of Q-indices and amplitude support the hypothesis that our primary QPD sample is composed of two distinct populations of variables with rotation and intermediate timescale, where the rotation timescale QPDs have lower amplitude on average (0.29 mag) and are more periodic (lower Q), while the intermediate timescale QPDs tend to have larger amplitude on average (0.44 mag) and reduced periodicity in comparison. We note that this second population of intermediate-timescale QPDs is also present in the ZTF sample, which has a timespan close to Gaia’s, suggesting that this second type of QPDs is absent from high-cadence surveys due to their limited total timespan.
@@
While QPDs at rotational timescales can be associated with AA Tau type variability, QPD with intermediate timescales could be related to occultation by disk-material located farther away than the inner disk. One possible prototype for this variability behaviour could be the young star RW Aur (Cabrit et al., 2006; Facchini et al., 2016). A proposed phenomenological explanation for the fading of RW Aur would be a tidally truncated disk in which a flyby removed the outer part of the disk, leaving a tidally disrupted ’arm’ that induces occultations of stellar light when observed from moderate viewing angles (e.g. Rodriguez et al., 2013). We note that although RW Aur is best known for its 2 mag deep fading events in 2010 and 2014, lower amplitude variability could be expected under the same phenomenology (e.g Fischer et al., 2023). A different possible phenomenon could be disk instabilities as a trigger of dense disk winds responsible for enhancing line-of-sight extinction above the disk mid-plane, which, due to their radial location, would cause rotational modulation at longer timescales (e.g. Bouvier et al., 2013). Further constraints on the colour behaviour during dipping features could help further probing of QPD.
@@ 6.2.4 Aperiodic Dippers
APD-variables (see Fig. 18) have the highest inner disk and CTTS frequencies (56.94 and 39.64%), pointing to an association of their observed variability with the star-disk interaction process. Previous studies have also linked APDs to the physical phenomenology of AA Tau. However, while QPDs require star-disk systems that undergo stable accretion to allow dust to be lifted above the disk midplane long enough for the occultations to be observed as quasiperiodic, in APDs, an unstable accretion regime will instead result in stochastically distributed aperiodic dippers through the light curve (Kulkarni & Romanova, 2008; McGinnis et al., 2015). Nevertheless, material occulting the star is still expected to be at around the co-rotation radius, thus aperiodic occultation features are still expected to have durations around rotational timescales. In contrast, Tables 3 and Fig. 14 reveal that APDs show intermediate timescales 2.26 times more often than rotational ones.
@@
We note that a meaningful interpretation of aperiodic timescales can be tricky. Fig. 14 shows that APDs from high-cadence surveys have timescales between a week and a month, with these surveys being limited to a time span of 40-80 days. May their APDs be in reality an undersampled quasiperiodic phenomenon occurring over timescales longer than their timespan, these will be detected as quasiperiodic by surveys like Gaia or ZTF, which explains the population of intermediate timescales among our QPDs, and the reduced population of rotation timescales in our primary APDs sample. Second, most aperiodic light curves had their timescale determined by the SF in Sect. 4.2.2. SF aperiodic timescales measure the time it takes for the dominant variability feature in the light curve to reach its maximum amplitude, thus reflecting the duration of the variability feature, rather than the frequency of occurrence of variability over the span of observations. For example, the bottom panel of Fig. 18 shows an APD light curve potentially linked to the AA Tau phenomena, where the 9.3 d derived aperiodic timescale reflects the time to reach the maximum amplitude (0.24 mag) in the main dip visible in the light curve. In the top panel of the same figure, we show a second example where the timescale of 147.7 days reflects half of the duration of a main fading event with amplitude 0.53 mag. Combined with the large amplitudes found among our primary sample of APDs, the large disk and CTTS frequencies and large aperiodic timescales support the suggestions that beyond AA Tau-like light curves, APDs at intermediate timescales may be a population of moderate-amplitude RW Aur fading light curves (e.g. Petrov et al., 2015).

### v2

@@ 6.1 Incidence of morphological classes across surveys
We also point to survey-to-survey differences in the time span of observations. High-cadence surveys cover between 40 and 80 days of observations, while ZTF surveys cover 800–1 500 days, and GDR3 750–1 000 days. These differences imply that quasi-periodic phenomena at intermediate timescales may be perceived as aperiodic by studies with reduced time span. Moreover, L variables are defined as sources with timescales larger than 20-40 d in high-cadence studies, 100-250 d in ZTF, and we used 350-500 d in the present study. L variable fractions are thus fully survey-dependent. Even if removed from our class incidence comparison, their impact on the definition of other morphological classes is still worth mentioning. First, L variables could be operationally defined as aperiodic, since their time scale was not resolvable over the duration of a survey. Although their lack of reliable timescale implies that their Q indices cannot be estimated, accounting for their M indices would increase class frequencies by 3.8% for B, 8.6% for S, and 5.3% for APDs - thus bringing our class incidence for S and APD variables to levels similar to those of ZTF studies. Second, the timescale limit for the L variable definition indirectly determines how many folds are used in the waveform estimation required to derive the Q index. With our threshold of half of the total duration of observations, only one fold is included for sources with the longest timescales, which may lead to over-fitting of estimated waveforms, resulting in deflated Q indices, and causing aperiodic phenomena to be miss-detected as quasi-periodic.
@@
Some previous studies attributed multiple morphological classes to multi-mode variables (e.g. Cody_2018; Cody_2022; Hillenbrand2022) also including an additional class, multi-periodic, to account for variables with multiple periods. In Table 2, multi-mode variables were represented by their primary morphological class. However, with our focus on the dominant variability mode, we acknowledge that undetected multi-mode variables in our sample may display inflated Q indices. Beyond multiple variability mechanisms occurring simultaneously, the longer time span of surveys like Gaia and ZTF may also facilitate the observation of sources with alternating dominant variability phenomena. For example, this can result in symmetric M indices derived from light curves that shift from bursting to dipping behaviour (e.g., Fig. 22).
@@
Our total fraction of dipping variables (QPD and APD) matches that reported in ZTF studies, but is lower than in high-cadence surveys. Nevertheless, as already noted, our APD fractions are underestimated due to our definition of L variables. Furthermore, in Sects. 6.2.3 and 6.2.4 we discuss evidence that our sample of dippers includes a different underlying population of YSOs in comparison to the high-cadence surveys. Similarly, B variables also include different variability mechanisms, and in Sect. 6.2.5 we discuss evidence that the Gaia sampling yields the detection of a distinct population of bursting variables in comparison with other surveys.
6.2 General properties of the morphological classes
@@ 6.2.1 Strictly periodic
Strictly periodic variables (P, Fig. 17) are usually associated with rotational modulation by magnetically induced cool spots at the stellar surface in sources viewed close to edge-on (; Cody_2022). Fig. 16 shows that the timescale distribution for P sources is consistent across studies and dominated by rotational timescales. The variability amplitude in our primary sample (blue) is higher than that reported in high-cadence surveys (red) but slightly lower than in ZTF studies (orange); the secondary sample (grey) exhibits amplitude levels comparable to those of high-cadence studies. Differences in amplitude distributions can be attributed to selection effects in different surveys. Typically, cool spots have lifetimes ranging from 1 to 50 rotation cycles (Basri2022ApJ...924...31B). Higher-cadence surveys, which usually span 40 to 80 d, provide excellent phase coverage before significant changes in spot coverage occur. In Gaia, sparse sampling and long-term coverage implies that hundreds of rotation cycles take place before achieving sufficient phase coverage, thus biasing the sample of P sources towards sources with larger and more stable spot coverage (and large amplitudes). In contrast, the higher noise level of ground-based observations allows only the higher-amplitude variables to be detected. As a result, Fig. 16 shows that the ZTF distributions of all classes are shifted toward higher amplitudes compared to other surveys.
The properties of our P variables support the hypothesis that their variability is primarily driven by rotational modulation linked to relatively stable cool spots. Their average amplitude is among the lowest of all morphological classes, and the ratio between rotation and intermediate timescales (Table 1) shows that this sample is dominated by rotational modulation. P variables are 82.95% composed of YSOs without evidence of a thick disc or envelope, and 90.2% composed of K- and M-type stars. Furthermore, 86.27% of P sources with H information are WTTS.
@@ 6.2.2 Quasi-periodic symmetric
Multiple physical origins have been proposed to explain QPS variables (Fig. 18), which are the most common in our entire sample. Although generally linked to rotational modulation of stellar spots, high-amplitude QPS have been tentatively associated with variability caused by large hot spots induced by accretion, in combination with underlying lower-amplitude accretion variation (Venuti2015A&A...581A..66V; Cody_2018), while low-amplitude QPS are thought to originate from changes of cool spot coverage due to magnetic activity (Cody_2022). Considering that cool spots are the only variability driver expected in discless sources and that disc-bearing and actively accreting sources are likely to exhibit accretion-induced hot spots, this view is generally supported by our sample. We find that the average distribution of amplitudes is lower for discless and thin-disc QPS (0.11 mag) compared to sources with envelope or thick disc (0.30). Similar differences are found when comparing the amplitudes of the WTTS QPS (0.11) with the CTTS QPS (0.36). The same trend is verified by comparing our primary and secondary samples, where the CTTS frequency varies from 36.6% to 6.1%. The link of QPS with rotational modulation by spots is further reinforced by the observation that M-type QPS have an excess of short timescales, with 60% of them displaying timescales of days compared to for other spectral types. This is in accordance with the expectation that M-type stars tend to be faster rotators than their more massive counterparts (e.g. 2007ApJ...671..605C).
Our QPS-variables include a population of sources with intermediate timescales, which has not been found in previous studies. Fig. 16 shows that for the primary sample, the timescale distribution is bimodal, with 1.3 times as many rotation timescales as intermediate ones. A comparison of their Q index distributions reveals significant differences (with KS test support), where the rotational-timescale QPS are shifted toward higher periodicity with an average Q of 0.34 compared to 0.44 for intermediate-timescale QPS.
@@
Visual inspection of intermediate-timescale QPS light curves revealed that many of these sources have limited phase coverage when folded to their characteristic timescale. This limitation is inherent in surveys with sparse sampling as Gaia, where some phase-folded light curves include large gaps with no data. These gaps, in turn, bring us to the limits of the method for waveform estimation, causing waveforms to be over-fit to only a fraction of the phase, artificially shifting Q indices towards lower values and causing misdetection of aperiodic sources as quasi-periodic. With the Gaia scanning law, the timescales that yield gaps in phase coverage vary with sky position, making modelling and masking these timescales from our analysis impractical. To mitigate this issue, future studies should consider the complementation of Q-M analysis with an additional variability metric sensitive to the degree of phase coverage as part of the procedure for the derivation of the Q index (e.g. Saunders2006AN....327..783S). We also considered adjusting the quasi-periodic/aperiodic Q index threshold to mitigate this effect, but verified that a lower Q index limit would result in reliable quasi-periodic light curves with rotation timescales being misclassified as aperiodic.
@@ 6.2.3 Quasi-periodic dippers
The phenomenology AA Tauri variables (Bouvier1999A&A...349..619B) is often linked to QPD light curves (see Fig. 19), where a misalignment between the rotation and magnetic field axes yields the formation of wraps, with disc material lifted above the disc mid-plane and close to the co-rotation radius, introducing quasi-periodic occultations at rotational timescales in systems observed close to edge-on (Romanova2013MNRAS.430..699R; McGinnis2015A&A...577A..11M). Among the primary sample, 62.09% of QPDs have an inner disc, and 40.60% have evidence of ongoing accretion (CTTS classification), supporting the association of this class with physical phenomenology related to the presence of a disc and ongoing accretion. However, a direct association of these with the AA Tau prototype is hampered by two factors.
@@
Firstly, a fraction of our low-amplitude QPDs may be misclassified QPS. About half of our QPDs are in the secondary sample, which shows a much lower disc (13.71%) and CTTS (4.47%) frequencies, with both Q and M distributions shifted towards more symmetric M and more periodic Q indices compared to the primary sample. Fig. 16 shows that these have systematically lower amplitudes than QPDs from high-cadence surveys. The M index distribution for the secondary QPD sample peaks near the threshold for dipper selection, with 90% of the M indices below the median value for the primary QPD sample. Additionally, although statistically different in shape, the distributions of amplitudes and Q index for the secondary QPD and secondary QPS samples have similar averages and standard deviations, as well as similar disc and CTTS frequency. Finally, visual examination of the light curves also suggested that a higher M index threshold might be necessary to enhance the purity of the dipper variable samples, where a reviewed threshold would identify a large fraction of secondary QPD as QPS.
@@
Secondly, because the observation time span of Gaia is much longer, QPDs may be sensitive to other dimming variables beyond AA Tau. Fig. 16 reveals a bimodal distribution for QPDs, with of QPDs showing intermediate timescales. This contradicts the predictions of the AA Tau phenomenology, where the occulting material is anticipated to be near the co-rotation radius. The primary sample of QPDs with intermediate timescales shows high disc and CTTS frequencies (65.09, and 47.35%), suggesting a physical origin also associated with the presence of a disc and ongoing accretion. Statistical tests for distribution shape, mean and variance differences for the distributions of Q indices and amplitude support the hypothesis that our primary QPD sample is composed of two distinct populations of variables with rotation and intermediate timescale, where the rotation timescale QPDs have lower amplitude on average (0.29 mag) and are more periodic (lower Q), while the intermediate timescale QPDs tend to have larger amplitude on average (0.44 mag) and reduced periodicity in comparison. We note that this second population of intermediate-timescale QPDs is also present in the ZTF sample, which has a time span close to that of Gaia, suggesting that this second type of QPDs is absent from high-cadence surveys due to their limited total time span.
@@
While QPDs at rotational timescales can be associated with AA Tau type variability, those with intermediate timescales could be related to occultation by disc material located farther away than the inner disc. One possible prototype for this variability behaviour could be the young star RW Aur (Cabrit2006A&A...452..897C; Facchini2016A&A...596A..38F). A proposed phenomenological explanation for the fading of RW Aur would be a tidally truncated disc in which a flyby removed the outer part of the disc, leaving a tidally disrupted ’arm’ that induces occultations of stellar light when observed from moderate viewing angles (e.g. Rodriguez2013AJ....146..112R). We note that although RW Aur is best known for its 2 mag deep fading events in 2010 and 2014, lower amplitude variability could be expected under the same phenomenology (e.g Fischer23). A different possible phenomenon could be disc instabilities as a trigger of dense disc winds responsible for enhancing line-of-sight extinction above the disc mid-plane, which, due to their radial location, would cause orbital modulation at longer timescales (e.g. Bouvier2013A&A...557A..77B). Further constraints on the colour behaviour during dipping features could help further probe QPDs.
@@ 6.2.4 Aperiodic dippers
With the highest inner disc and CTTS frequencies (56.94 and 39.64%), APD variability (Fig. 20) is likely associated with the star-disc interaction process. Previous studies have also linked APDs to the physical phenomenology of AA Tau. However, while QPDs require star-disc systems that undergo stable accretion to allow dust to be lifted above the disc midplane long enough for the occultations to be observed as quasi-periodic, in APDs, an unstable accretion regime will instead result in stochastically distributed aperiodic dippers through the light curve (KulkarniRomanova2008MNRAS.386..673K; McGinnis2015A&A...577A..11M). Nevertheless, material occulting the star is still expected to be at around the co-rotation radius, thus aperiodic occultation features are still expected to have durations around rotational timescales. In contrast, Tables 1 and Fig. 16 reveal that APDs show intermediate timescales 2.26 times more often than rotational ones.
@@
We note that a meaningful interpretation of aperiodic timescales can be tricky. Fig. 16 shows that APDs from high-cadence surveys have timescales between a week and a month, with these surveys being limited to a time span of 40-80 days. If their APDs are in reality an undersampled quasi-periodic phenomenon occurring over timescales longer than their time span, these will be detected as quasi-periodic by surveys like Gaia or ZTF, which explains the population of intermediate timescales among our QPDs, and the reduced population of rotation timescales in our primary APDs sample. Second, most aperiodic light curves had their timescale determined by the SF in Sect. 4.2.2. SF aperiodic timescales measure the time it takes for the dominant variability feature in the light curve to reach its maximum amplitude, thus reflecting the duration of the variability feature, rather than the frequency of occurrence of variability over the span of observations. For example, the bottom panel of Fig. 20 shows an APD light curve potentially linked to the AA Tau phenomena, where the 9.3 d derived aperiodic timescale reflects the time to reach the maximum amplitude (0.24 mag) in the main dip visible in the light curve. In the top panel of the same figure, we show a second example where the timescale of 147.7 days reflects half of the duration of a main fading event with amplitude 0.53 mag. Combined with the large amplitudes found among our primary sample of APDs, the large disc and CTTS frequencies and large aperiodic timescales support the suggestions that beyond AA Tau-like light curves, APDs at intermediate timescales may be a population of moderate-amplitude RW Aur fading light curves (e.g. Petrov2015).
With 8.3% of APDs having estimated spectral types between AB and F and 87% of these having a disc, a possible explanation for these earlier type APDs could be the phenomenology behind UX Ori variables (Natta1997ApJ...491..885N). These YSOs have their variability attributed to large-scale disc perturbations causing a puffing-up of disc material close to the dust sublimation radius (2022MNRAS.512.3098S), which, when observed almost edge-on, results in sporadic and large amplitude (1–2 mag) fades. Nevertheless, we verified that although the distributions of amplitudes of earlier and later type APDs are statistically different, the former have much lower typical amplitudes ( mag) compared to the latter ( mag).